# IUST_PersonReId: A New Domain in Person Re-Identification Datasets

### v1

@@ Abstract
Person re-identification (ReID) models often struggle to generalize across diverse cultural contexts, particularly in Islamic regions like Iran, where modest clothing styles are prevalent. Existing datasets predominantly feature Western and East Asian fashion, limiting their applicability in these settings. To address this gap, we introduce IUST_PersonReId, a dataset designed to reflect the unique challenges of ReID in new cultural environments, emphasizing modest attire and diverse scenarios from Iran, including markets, campuses, and mosques.
Experiments on IUST_PersonReId with state-of-the-art models, such as Solider and CLIP-ReID, reveal significant performance drops compared to benchmarks like Market1501 and MSMT17, highlighting the challenges posed by occlusion and limited distinctive features. Sequence-based evaluations show improvements by leveraging temporal context, emphasizing the dataset’s potential for advancing culturally sensitive and robust ReID systems. IUST_PersonReId offers a critical resource for addressing fairness and bias in ReID research globally. The dataset is publicly available at https://computervisioniust.github.io/IUST_PersonReId/.
@@
Person Re-identification , IUST_PersonReId , Demographic Bias , Dataset , Cultural Clothing
@@ 1 Introduction
Person ReIDentification (ReID) is a crucial task in computer vision, aimed at identifying and matching individuals across images captured by different cameras at various times and locations. The primary motivation for ReID arises from its vital applications in surveillance, security, and urban management. Surveillance cameras, often deployed across vast areas, generate massive amounts of footage. Manually sifting through this data to track individuals is impractical. Thus, automated ReID systems provide a scalable and efficient solution for matching identities across non-overlapping camera views, enabling more effective monitoring and management of public spaces.
@@
The performance of AI models, heavily depends on the quality and diversity of their training datasets, and ReID models are no exception. Developing robust models requires datasets that capture a wide range of variations in lighting, viewpoints, quality, and human appearances. Among these factors, clothing is a significant determinant of human appearance, varying dramatically between different cultures. For instance, women’s clothing in countries where hijabs are commonly worn, such as Iran, differs markedly from clothing in other regions. Incorporating cultural clothing diversity in ReID datasets is essential. These datasets not only improve model accuracy and reliability but also ensure fairness by preventing demographic bias in real-world applications [1]. Without sufficient representation, models may perform poorly on underrepresented groups, leading to biased and unfair outcomes.
@@
In recent years, several person ReID datasets have been developed, each addressing unique challenges. The CUHK03 [2] dataset captures 1,364 pedestrians across six surveillance cameras, introducing cross-view complexities with varying pedestrian movements. The DukeMTMC-reID [3, 4] dataset, derived from DukeMTMC, includes 1,852 identities across eight cameras, tackling real-world challenges like occlusions, illumination changes, detection errors, and low-resolution imagery. Market-1501 [5] provides over 32,000 annotated bounding boxes from six cameras, featuring distractor images and intra-class variations in a supermarket setting. MSMT17 [6] spans 15 cameras and captures 4,101 identities under varying lighting, weather, and time conditions, testing model robustness in both indoor and outdoor environments.
@@
Unlabeled datasets have gained significant attention for pre-training deep models. These datasets can easily scale, reduce data collection costs, and are particularly valuable for learning robust feature representations. LUPerson [7], for instance, comprises 4 million images from 200,000 identities, predominantly derived from studio-quality videos on YouTube, rather than real-world surveillance conditions. This dataset spans diverse scenes such as streets, campuses, and sports fields, enabling large-scale pre-training for ReID. LUPerson-NL [8], an extension of LUPerson with 10 million images and 430,000 identities, introduces noisy labels through automated tracking, offering greater diversity in person representations.
@@
To tackle these challenges, we present the IUST_PersonReId dataset, which encompasses unique cultural and environmental conditions. This dataset is designed to fill the existing gap and provide a robust resource for training and evaluating ReID models in contexts where modest clothing is prevalent. We provide an extensive evaluation procedure to validate our dataset and its significance for this domain. We show that the observed performance decrease due to cultural clothing differences is real, and the dataset effectively addresses this issue. Our evaluations demonstrate significant improvements in the new domain, reflected by an increase in rank-1 accuracy by about 40%.
The structure of the paper is as follows: Section 2 reviews related works. Section 3 describes data collection, annotation, and dataset composition. Section 4 discusses model performance, the impact of cultural clothing, and key insights. Finally, Section 5 provides the conclusions and potential directions for future work.
@@ 2.1 Supervised Datasets
The CUHK03 dataset [2], created by the Chinese University of Hong Kong, consists of images and videos captured on the university’s campus using two pairs of cameras. The dataset represents East Asian modern clothing culture and includes both manually labeled and automatically detected bounding boxes, with the latter generated by the DPM algorithm [9]. The Market-1501 dataset [5] is a widely used resource collected at the Tsinghua University campus supermarket with up to six cameras, capturing East Asian modern clothing culture and including detections from the DPM algorithm [9]. An extension of this dataset, MARS [10], was also collected on the Tsinghua University campus using up to six cameras. It encompasses clothing styles same as Market-1501, with detections provided by both the DPM and GMMCP [11] algorithms.
@@
The DESTRE-P dataset [19] was collected in two different countries, India and Portugal, and represents the clothing culture of both regions in an academic environment. Unlike other datasets, it is not limited to a single location or country. The LaST [20] dataset created by collecting 2k movies, covering more than 8 countries from Asia to Europe. This dataset focus on enlarging spatial scope and time span of pedestrian activities by using movies sources. This dataset covers more diverse clothing culture in comparison of previous datasets by covering multiple countries.
@@
The MSMT17 dataset [6] was collected on a university campus. However, it lacks specific information regarding the location and the clothing culture represented within the dataset. The LPW dataset [21] comprises 2,731 identities, each captured by up to four cameras. However, there is no information available about the location where the dataset was collected. The MALS dataset [22] is a recent addition to the field, designed to address the challenges of text-based person retrieval through a large-scale, multi-attribute, and language search benchmark. The dataset is generated to simulate diverse pedestrian appearances, environments, and multi-modal queries. It includes over 1.5 million synthetic images and associated textual descriptions, making it one of the largest datasets in this domain. This dataset does not adhere to the clothing styles of any specific culture or country, as it was entirely generated using synthetic methods.
@@ 2.2 Unsupervised Datasets
Unsupervised datasets have gained significant attention in recent years due to their ability to scale easily and reduce the costs associated with data collection and labeling. The LUPerson dataset [7] comprises over 4 million pedestrian images, estimated to include over 200k identities, sourced from a diverse range of environments, aimed at improving unsupervised person re-identification (ReID) models through the provision of extensive unlabeled data. The dataset was curated by crawling over 70,000 street-view videos from YouTube, using search queries such as “city name + street view (or scene)”. To ensure broad diversity, the dataset includes videos from the top 100 largest cities globally, predominantly located in Western countries and East Asia, thereby reflecting a wide array of clothing cultures. Building upon LUPerson, the LUPerson-NL dataset [8] introduces noisy labels to simulate real-world labeling errors, which is crucial for the development and evaluation of ReID algorithms capable of handling label noise. Each tracklet generated by the tracking algorithm is considered a distinct identity and referred to as a noisy label. This extension is designed to ensure that ReID models can maintain high performance even when confronted with inaccuracies in the training data. Table 1 provides a summary of the datasets mentioned above.
@@
3 Data Collection & Characteristics
Our dataset was collected with a specific focus on the clothing styles prevalent in Islamic countries, particularly emphasizing women who wear hijab.
3.1 Collection
@@ 3.3 Annotation Process
In this phase the raw videos split into 30 minutes partitions. After that we use pedestrian tracking algorithms to pre-annotate each of the partitions. Here we use multiple tracking algorithms suitable for each environment cause each of the algorithms surpasses other algorithms in specific environments and we use them all. We use YOLOv5 [23], YOLOv8 [23] with bytetrack [24] tracker, FairMOT [25] and YOLOE [26] models which have builtin trackers. After auto annotation step, we conduct a filtering procedure to make supervised annotation process much easier. We remove bounding boxes which appeared in less than 5 frames because during the annotation these bounding boxes appear for a very short time and disappear quickly and can’t be easily seen by annotator to validate or edit it. We use CVAT [27] application for make corrections in automated annotations and make it available online to all of annotators to make them work remotely. To help them download the frames much faster we conduct resizing the videos with respect to their file size with using compression strategies available in this platform.
@@
The generation of supervised data for the re-identification task involves annotating data to match the same individuals across different camera views (cross-camera re-identification) and within the same camera as they exit and re-enter the scene (self re-identification). Although tools such as ROSE ReID [28] and the Semi-Automatic Data Annotation Tool [29] exist, they are not publicly accessible. Consequently, we developed our own annotation tool for the re-identification phase, named CVLab-ReID-Tool  [^1] https://github.com/ComputerVisionIUST/CVLab-ReId-Tool, which we have made publicly available. An example of the designed app can be seen in Figure 2.
@@ 3.4 Sampling Procedure
Frames are sampled from each annotated video at intervals of 250 ms. Since the raw videos have different frame rates (FPS), this interval corresponds to $0.25\times\mathrm{FPS}$, giving the frame number interval equivalent to 250 ms. This approach ensures uniform sampling across videos with varying FPS. Additionally, we limit the number of frames for each individual to a maximum of 50, which helps reduce redundancy and dataset size while maintaining sufficient diversity. For individuals with fewer than 50 frames, all available frames are included.
@@
For individuals with more than 50 sampled frames, we employ a quality assessment algorithm to select the best-quality frames. The BRISQUE [30] algorithm is used to assess the visual quality of each frame. Using a sliding window approach, we evaluate consecutive subsets of frames and select the subset with the highest aggregate quality score. This ensures the dataset prioritizes clear and visually informative frames, which are essential for robust model training and evaluation.
@@ 4 Results
This section presents the evaluation of our proposed dataset for person re-identification (ReID) tasks. We outline the dataset preparation process, describe the training and evaluation methodologies, and provide performance metrics to assess the effectiveness of the models and dataset. Detailed results include baseline evaluations, sequence-based re-identification, and various ablation studies such as cross-dataset performance, gender-based analysis, and visibility-based evaluations.
@@ 4.1 Dataset Preparation
4.2 Models and Training
4.3 Evaluation Metrics
Cumulative Match Characteristic (CMC)
The Cumulative Match Characteristic (CMC) curve is a key metric for assessing person re-identification (ReID) performance. It calculates the probability of a correct match appearing within the top-kk ranks for a given query. We report the Rank-1, Rank-5, and Rank-10 accuracies as standard measures of model performance [34].
Mean Average Precision (mAP)
@@ 4.6 Ablation Studies
Cross-Dataset Performance Comparison
@@
Table 5: Cross-dataset performance comparison (%) for Solider. The numbers in parentheses represent the amount of performance drop relative to the diagonal elements. Rows represent the training dataset, while columns represent the testing dataset.
Train / Test Market1501 Duke MSMT17 IUST_PersonReId
Market1501 93.10 52.90 ($\downarrow 40.20$) 16.95 ($\downarrow 76.15$) 5.01 ($\downarrow 88.09$)
Duke 14.21 ($\downarrow 78.89$) 52.89 1.12 ($\downarrow 51.77$) 0.71 ($\downarrow 52.18$)
MSMT17 50.05 ($\downarrow 15.31$) 51.71 ($\downarrow 13.65$) 65.36 8.12 ($\downarrow 57.24$)
IUST_PersonReId 30.15 ($\downarrow 12.20$) 31.06 ($\downarrow 11.29$) 9.30 ($\downarrow 33.05$) 42.35
Table 6: Cross-dataset performance comparison (%) for CLIP-ReID. The numbers in parentheses represent the amount of performance drop relative to the diagonal elements. Rows represent the training dataset, while columns represent the testing dataset.
Train / Test Market1501 Duke MSMT17 IUST_PersonReId
Market1501 89.67 51.09 ($\downarrow 38.58$) 23.52 ($\downarrow 66.15$) 6.69 ($\downarrow 82.98$)
Duke 41.73 ($\downarrow 41.04$) 82.77 21.72 ($\downarrow 61.05$) 5.77 ($\downarrow 77.00$)
MSMT17 46.12 ($\downarrow 27.20$) 57.60 ($\downarrow 15.72$) 73.32 13.24 ($\downarrow 60.08$)
IUST_PersonReId 39.63 ($\downarrow 11.95$) 42.92 ($\downarrow 8.66$) 19.24 ($\downarrow 32.34$) 51.58
@@
Gender-based Query and Gallery Setup

### v2

# A Culturally-Aware Benchmark for Person Re-Identification in Modest Attire

@@ Abstract
Person Re-Identification (ReID) is a fundamental task in computer vision with critical applications in surveillance and security. Despite progress in recent years, most existing ReID models often struggle to generalize across diverse cultural contexts, particularly in Islamic regions like Iran, where modest clothing styles are prevalent. Existing datasets predominantly feature Western and East Asian fashion, limiting their applicability in these settings. To address this gap, we introduce Iran University of Science and Technology Person Re-Identification (IUST_PersonReId), a dataset designed to reflect the unique challenges of ReID in new cultural environments, emphasizing modest attire and diverse scenarios from Iran, including markets, campuses, and mosques.
Experiments on IUST_PersonReId with state-of-the-art models, such as Semantic Controllable Self-supervised Learning (SOLIDER) and Contrastive Language-Image Pretraining Re-Identification (CLIP-ReID), reveal significant performance drops compared to benchmarks like Market1501 and Multi-Scene MultiTime (MSMT17), specifically, SOLIDER shows a drop of 50.75% and 23.01% Mean Average Precision (mAP) compared to Market1501 and MSMT17 respectively, while CLIP-ReID exhibits a drop of 38.09% and 21.74% mAP, highlighting the challenges posed by occlusion and limited distinctive features. Sequence-based evaluations show improvements by leveraging temporal context, emphasizing the dataset’s potential for advancing culturally sensitive and robust ReID systems. IUST_PersonReId offers a critical resource for addressing fairness and bias in ReID research globally.
@@
Person Re-identification , Dataset , Demographic Bias , Cultural Clothing
@@ 1 Introduction
Person Re-Identification (ReID) is a crucial task in computer vision, aimed at identifying and matching individuals across images captured by different cameras at various times and locations. The primary motivation for ReID arises from its vital applications in surveillance, security, and urban management. Surveillance cameras, often deployed across vast areas, generate massive amounts of footage. Manually sifting through this data to track individuals is impractical. Thus, automated ReID systems provide a scalable and efficient solution for matching identities across non-overlapping camera views, enabling more effective monitoring and management of public spaces.
@@
The performance of Artificial Intelligence (AI) models, heavily depends on the quality and diversity of their training datasets, and ReID models are no exception. Developing robust models requires datasets that capture a wide range of variations in lighting, viewpoints, quality, and human appearances. Among these factors, clothing is a significant determinant of human appearance, varying dramatically between different cultures. For instance, women’s clothing in countries where hijabs are commonly worn, such as Iran, differs markedly from clothing in other regions. Incorporating cultural clothing diversity in ReID datasets is essential. These datasets not only improve model accuracy and reliability but also ensure fairness by preventing demographic bias in real-world applications [1]. Without sufficient representation, models may perform poorly on underrepresented groups, leading to biased and unfair outcomes.
@@
In recent years, several person ReID datasets have been developed, each addressing unique challenges. The Chinese University of Hong Kong (CUHK03) [2] dataset captures 1,364 pedestrians across six surveillance cameras, introducing cross-view complexities with varying pedestrian movements. The Duke Multi-Target, Multi-Camera Re-Identification (DukeMTMC-reID) [3, 4] dataset, derived from DukeMTMC, includes 1,852 identities across eight cameras, tackling real-world challenges like occlusions, illumination changes, detection errors, and low-resolution imagery. Market-1501 [5] provides over 32,000 annotated bounding boxes from six cameras, featuring distractor images and intra-class variations in a supermarket setting. MSMT17 [6] spans 15 cameras and captures 4,101 identities under varying lighting, weather, and time conditions, testing model robustness in both indoor and outdoor environments.
@@
Unlabeled datasets have gained significant attention for pre-training deep models. These datasets can easily scale, reduce data collection costs, and are particularly valuable for learning robust feature representations. Large-scale Unlabeled Person Re-Identification (LUPerson) [7], for instance, comprises 4 million images from 200,000 identities, predominantly derived from studio-quality videos on YouTube, rather than real-world surveillance conditions. This dataset spans diverse scenes such as streets, campuses, and sports fields, enabling large-scale pre-training for ReID. LUPerson-NL [8], an extension of LUPerson with 10 million images and 430,000 identities, introduces noisy labels through automated tracking, offering greater diversity in person representations.
@@
To tackle these challenges, we present the IUST_PersonReId dataset, which encompasses unique cultural and environmental conditions. This dataset is designed to fill the existing gap and provide a robust resource for training and evaluating ReID models in contexts where modest clothing is prevalent. We provide an extensive evaluation procedure to validate our dataset and its significance for this domain. We show that the observed performance decrease due to cultural clothing differences is real, and the dataset effectively addresses this issue. Our evaluations demonstrate significant improvements in the new domain, reflected by an increase in rank-1 accuracy by about 40%. The dataset is publicly available at https://computervisioniust.github.io/IUST_PersonReId/.
The structure of the paper is as follows: Section 2 reviews related works. Section 3 describes methodology of data collection, annotation, and dataset composition. Section 4 discusses model performance, the impact of cultural clothing, and key insights. Finally, Section 5 provides the conclusions and potential directions for future work.
@@ 2.1 Supervised Datasets
The CUHK03 dataset [2], created by the Chinese University of Hong Kong, consists of images and videos captured on the university’s campus using two pairs of cameras. The dataset represents East Asian modern clothing culture and includes both manually labeled and automatically detected bounding boxes, with the latter generated by the Deformable Parts Model (DPM) algorithm [9]. The Market-1501 dataset [5] is a widely used resource collected at the Tsinghua University campus supermarket with up to six cameras, capturing East Asian modern clothing culture and including detections from the DPM algorithm [9]. An extension of this dataset, Motion Analysis and Re-identification Set (MARS) [10], was also collected on the Tsinghua University campus using up to six cameras. It encompasses clothing styles same as Market-1501, with detections provided by both the DPM and Generalized Maximum Multi-Clique Problem (GMMCP) [11] algorithms.
@@
The Pedestrian Detection, Tracking, and Re-Identification (DESTRE-P) dataset [19] was collected in two different countries, India and Portugal, and represents the clothing culture of both regions in an academic environment. Unlike other datasets, it is not limited to a single location or country. The Large-scale Spatio-Temporal Person Re-Identification (LaST) [20] dataset created by collecting 2k movies, covering more than 8 countries from Asia to Europe. This dataset focus on enlarging spatial scope and time span of pedestrian activities by using movies sources. This dataset covers more diverse clothing culture in comparison of previous datasets by covering multiple countries.
@@
The MSMT17 dataset [6] was collected on a university campus. However, it lacks specific information regarding the location and the clothing culture represented within the dataset. The Labeled Pedestrian in the Wild (LPW) dataset [21] comprises 2,731 identities, each captured by up to four cameras. However, there is no information available about the location where the dataset was collected. The Multi-Attribute and Language Search (MALS) dataset [22] is a recent addition to the field, designed to address the challenges of text-based person retrieval through a large-scale, multi-attribute, and language search benchmark. The dataset is generated to simulate diverse pedestrian appearances, environments, and multi-modal queries. It includes over 1.5 million synthetic images and associated textual descriptions, making it one of the largest datasets in this domain. This dataset does not adhere to the clothing styles of any specific culture or country, as it was entirely generated using synthetic methods.
@@ 2.2 Unsupervised Datasets
Unsupervised datasets have gained significant attention in recent years due to their ability to scale easily and reduce the costs associated with data collection and labeling. The LUPerson dataset [7] comprises over 4 million pedestrian images, estimated to include over 200k identities, sourced from a diverse range of environments, aimed at improving unsupervised ReID models through the provision of extensive unlabeled data. The dataset was curated by crawling over 70,000 street-view videos from YouTube, using search queries such as “city name + street view (or scene)”. To ensure broad diversity, the dataset includes videos from the top 100 largest cities globally, predominantly located in Western countries and East Asia, thereby reflecting a wide array of clothing cultures. Building upon LUPerson, the LUPerson-NL dataset [8] introduces noisy labels to simulate real-world labeling errors, which is crucial for the development and evaluation of ReID algorithms capable of handling label noise. Each tracklet generated by the tracking algorithm is considered a distinct identity and referred to as a noisy label. This extension is designed to ensure that ReID models can maintain high performance even when confronted with inaccuracies in the training data. Table 1 provides a summary of the datasets mentioned above.
@@
3 Methodology
We followed a clear and structured process that involved collecting data from real-world settings, adding challenging conditions on purpose, carefully labeling the data, and using thoughtful sampling methods. The following subsections detail each of these components.
3.1 Data Collection
@@ 3.3 Annotation Process
In this phase the raw videos split into 30 minutes partitions. After that we use pedestrian tracking algorithms to pre-annotate each of the partitions. Here we use multiple tracking algorithms suitable for each environment cause each of the algorithms surpasses other algorithms in specific environments and we use them all. We use You Only Look Once version-5 (YOLOv5) [23], YOLOv8 [23] with bytetrack [24] tracker, Fair Multi Object Tracking (FairMOT) [25] and YOLOE [26] models which have builtin trackers. After auto annotation step, we conduct a filtering procedure to make supervised annotation process much easier. We remove bounding boxes which appeared in less than 5 frames because during the annotation these bounding boxes appear for a very short time and disappear quickly and can’t be easily seen by annotator to validate or edit it. We use Computer Vision Labelling Tool (CVAT) [27] application for make corrections in automated annotations and make it available online to all of annotators to make them work remotely. To help them download the frames much faster we conduct resizing the videos with respect to their file size with using compression strategies available in this platform.
@@
The generation of supervised data for the re-identification task involves annotating data to match the same individuals across different camera views (cross-camera re-identification) and within the same camera as they exit and re-enter the scene (self re-identification). Although tools such as Rapid-Rich Object Search (ROSE) ReID [28] and the Semi-Automatic Data Annotation Tool [29] exist, they are not publicly accessible. Consequently, we developed our own annotation tool for the re-identification phase, named Computer Vision Lab Re-Identification Tool (CVLab-ReID-Tool)  [^1] https://github.com/ComputerVisionIUST/CVLab-ReId-Tool, which we have made publicly available. An example of the designed app can be seen in Figure 2.
@@ 3.4 Sampling Procedure
Frames are sampled from each annotated video at intervals of 250 ms. Since the raw videos have different Frames Per Second (FPS) rates, this interval corresponds to $0.25\times\mathrm{FPS}$, giving the frame number interval equivalent to 250 ms. This approach ensures uniform sampling across videos with varying FPS. Additionally, we limit the number of frames for each individual to a maximum of 50, which helps reduce redundancy and dataset size while maintaining sufficient diversity. For individuals with fewer than 50 frames, all available frames are included.
@@
For individuals with more than 50 sampled frames, we employ a quality assessment algorithm to select the best-quality frames. The Blind Reference-less Image Spatial Quality Evaluator (BRISQUE) [30] algorithm is used to assess the visual quality of each frame. Using a sliding window approach, we evaluate consecutive subsets of frames and select the subset with the highest aggregate quality score. This ensures the dataset prioritizes clear and visually informative frames, which are essential for robust model training and evaluation.
@@ 3.5 Dataset Composition
3.6 Models and Training
3.7 Evaluation Metrics
Cumulative Match Characteristic (CMC)
The Cumulative Match Characteristic (CMC) curve is a key metric for assessing person re-identification (ReID) performance. It calculates the probability of a correct match appearing within the top-k ranks for a given query. We report the Rank-1, Rank-5, and Rank-10 accuracies as standard measures of model performance [34].
Mean Average Precision (mAP)
@@ 4 Results
This section presents the evaluation of our proposed dataset for person re-identification (ReID) tasks. We outline the dataset preparation process, describe the training and evaluation methodologies, and provide performance metrics to assess the effectiveness of the models and dataset. Detailed results include baseline evaluations, sequence-based re-identification, and various ablation studies such as cross-dataset performance, gender-based analysis, visibility-based evaluations, and facial feature analysis.
@@ 4.4 Ablation Studies
4.4.1 Cross-Dataset Performance Comparison
@@
Table 5: Cross-dataset performance comparison (mAP%). The numbers in parentheses represent the amount of performance drop relative to the diagonal element of the testing dataset. Rows represent the training dataset, while columns represent the testing dataset.
Model Train / Test Market1501 Duke MSMT17 DESTRE-P IUST_PersonReId
SOLIDER [31] Market1501 93.10 52.90 ($\uparrow 0.01$) 16.95 ($\downarrow 48.41$) 31.00 ($\downarrow 12.60$) 5.01 ($\downarrow 37.34$)
Duke 14.21 ($\downarrow 78.89$) 52.89 1.12 ($\downarrow 64.24$) 12.10 ($\downarrow 31.50$) 0.71 ($\downarrow 41.64$)
MSMT17 50.05 ($\downarrow 43.05$) 51.71 ($\downarrow 1.18$) 65.36 33.80 ($\downarrow 9.80$) 8.12 ($\downarrow 34.23$)
DESTRE-P 23.60 ($\downarrow 69.50$) 22.40 ($\downarrow 30.49$) 7.30 ($\downarrow 58.06$) 43.60 3.30 ($\downarrow 39.05$)
IUST_PersonReId 30.15 ($\downarrow 59.52$) 31.06 ($\downarrow 21.83$) 9.30 ($\downarrow 56.06$) 29.00 ($\downarrow 14.60$) 42.35
CLIP-ReID [32] Market1501 89.67 51.09 ($\downarrow 31.68$) 23.52 ($\downarrow 49.80$) 34.40 ($\downarrow 8.60$) 6.69 ($\downarrow 44.89$)
Duke 41.73 ($\downarrow 47.94$) 82.77 21.72 ($\downarrow 51.60$) 33.76 ($\downarrow 9.24$) 5.77 ($\downarrow 45.81$)
MSMT17 46.12 ($\downarrow 43.55$) 57.60 ($\downarrow 25.17$) 73.32 36.21 ($\downarrow 6.79$) 13.24 ($\downarrow 38.34$)
DESTRE-P 22.04 ($\downarrow 67.63$) 20.00 ($\downarrow 62.77$) 6.18 ($\downarrow 67.14$) 43.00 3.62 ($\downarrow 47.96$)
IUST_PersonReId 39.63 ($\downarrow 49.04$) 42.92 ($\downarrow 39.85$) 19.24 ($\downarrow 54.08$) 33.95 ($\downarrow 9.05$) 51.58
Table 6: Cross-dataset performance comparison (Rank-1%). The numbers in parentheses represent the amount of performance drop relative to the diagonal element of the testing dataset. Rows represent the training dataset, while columns represent the testing dataset.
Model Train / Test Market1501 Duke MSMT17 DESTRE-P IUST_PersonReId
SOLIDER [31] Market1501 96.55 70.19 ($\downarrow 0.18$) 34.71 ($\downarrow 45.64$) 73.90 ($\downarrow 10.40$) 8.80 ($\downarrow 42.28$)
Duke 32.51 ($\downarrow 64.04$) 70.37 3.41 ($\downarrow 76.94$) 43.80 ($\downarrow 40.50$) 1.30 ($\downarrow 49.78$)
MSMT17 74.82 ($\downarrow 21.73$) 69.30 ($\downarrow 1.07$) 80.35 77.80 ($\downarrow 6.50$) 12.39 ($\downarrow 38.69$)
DESTRE-P 50.40 ($\downarrow 45.15$) 41.20 ($\downarrow 29.17$) 23.80 ($\downarrow 56.55$) 84.30 6.10 ($\downarrow 45.31$)
IUST_PersonReId 55.73 ($\downarrow 40.82$) 49.82 ($\downarrow 20.75$) 20.93 ($\downarrow 59.42$) 71.20 ($\downarrow 13.10$) 51.08
CLIP-ReID [32] Market1501 95.22 69.25 ($\downarrow 21.32$) 50.13 ($\downarrow 38.69$) 73.86 ($\downarrow 13.07$) 12.11 ($\downarrow 49.47$)
Duke 65.41 ($\downarrow 29.81$) 90.57 48.06 ($\downarrow 40.76$) 77.12 ($\downarrow 9.81$) 10.64 ($\downarrow 50.94$)
MSMT17 65.32 ($\downarrow 29.90$) 74.15 ($\downarrow 16.42$) 88.82 80.39 ($\downarrow 6.54$) 19.05 ($\downarrow 42.36$)
DESTRE-P 47.98 ($\downarrow 47.24$) 40.98 ($\downarrow 49.59$) 19.99 ($\downarrow 68.83$) 86.93 6.44 ($\downarrow 44.97$)
IUST_PersonReId 63.24 ($\downarrow 32.17$) 61.94 ($\downarrow 28.63$) 41.97 ($\downarrow 46.85$) 78.43 ($\downarrow 8.50$) 61.41
@@
4.4.2 Gender-based Query and Gallery Setup